OpenStudy (yellowlegoguy99):

A group of friends were voting for their favorite song: two-fifths chose song A, one-sixth of the remaining friends chose song B, and the rest chose song C. If there were thirty friends in all, how many chose song C? ten twelve fifteen eighteen

OpenStudy (yellowlegoguy99):

@E.ali

OpenStudy (anonymous):

Look we have : 1/5+1/6 = Persons chosed K ?!

OpenStudy (anonymous):

and : 1/5+1/6=6+5/30=11/30

OpenStudy (anonymous):

@yellowlegoguy99

OpenStudy (yellowlegoguy99):

11/30 isnt a result...?

OpenStudy (yellowlegoguy99):

Hellooo?

OpenStudy (anonymous):

Look : now : 30/30-11/30=19/30=persons in C Now have : 19__________________x 30_________________30 x=19 Persons C = 19 Got it ?:)

OpenStudy (yellowlegoguy99):

Yup ty :D

OpenStudy (yellowlegoguy99):

Wait...But 19 isnt a result either

OpenStudy (anonymous):

What ???

OpenStudy (yellowlegoguy99):

Go up to the top and look at the questions and you'll see the results

OpenStudy (anonymous):

oh ! caz I had a fall ... it was 2/5 not 1/5 ....

OpenStudy (anonymous):

Now go the last way and get 2/5 to 1/5 .

OpenStudy (yellowlegoguy99):

I came up with 17.........

OpenStudy (anonymous):

Again : 2/5+1/6=12+5/30 17/30=A & B 30/30-17/30=13/30=C And : 17_________________x 30________________30 x=17=Persons in C . But it isnt in answers !!!??? My answer is right !:)Mow !

OpenStudy (anonymous):

Now not Mow !

OpenStudy (yellowlegoguy99):

What??????????????????????????????